What is the speed of sound in a helium gas at a pressure of 15000pa

1
Expert's answer
2022-04-14T11:18:42-0400

The speed of sound in a gas

"v_s=\sqrt{\gamma\frac{P}{\rho}}=\sqrt{\gamma\frac{RT}{M}}"

At "t=0^\circ\rm C"

"v_s=\sqrt{\frac{5}{3}*\frac{8.31*273}{4*10^{-3}}}=972\:\rm m/s"